- Although we measured ADC noise with combinations of four ADC boards and three AA chassis, no combination in which ADC noise was nominal level was found.
- Finally, ADC1 was replaced from S1402719 to S1402710.
- Strange noise still remains, but it's better combination than others for now.
From the experience on ITMX, and ITMY, glitches and strange structure on spectra don't come from only ADC or AA.
So we tried some combinations and compared spectra of ADC1 noise in each combination.
Used instruments are as follows.
- S1402719: Originally installed as ADC1 at ETMY
- S1402710: Originally installed at ITMY. Finally installed as ADC1 at ETMY.
- S2214047: Another ADC (old product lot; revB)
- S2113341: A new product lot (revC). Tested also on ITMY
- S1706962: Originally and finally installed as AA1 at ETMY
- S2113203: Another AA (no modification)
- S2213719: Modified jumper pins (See also klog#19452)
Because we didn't replace instruments in ADC0 and ADC2 today, only one combination was measured for ADC0 and ADC2.
All results are put as /users/DGS/measurements/ADC/K1EY1/*.xml.
replacement of AA chassis
In the original situation, strange peak was able to be seen on many channels in addition to low-frequency excesses as shown in Fig.1 ~ Fig.4.
These behavior was not removed by replacing AA chassis. But frequency of peaks seemed to change by AA chassis replacement.
Fig.5 ~ Fig.8 and Fig.9 ~ Fig.12 shows the spectra with spare AA and with AA on which jumper pins were modified, respectively.
References in all figures represent a spectra in original situation.
We cannot see benefits by modifying jumper pins. But so many circuits are installed in EYV racks.
So we may need to take care about jumper pins on all circuits in order to see some benefit.
replacement of ADC board
By replacing ADC board, strange peaks around 10Hz vanished in all other combinations.
Fig.13 and Fig.14 shows the CH0-8 of ADC1 with another spare ADC and with ADC used in ITMY.
On the other hand, low-frequency excesses remained and channel number of contaminated channel depends on ADC boards we installed.
Unfortunately, there was no combination in which all 32 channels became healthy.
Finally, we installed ADC board which had been used in ITMY and the original AA chassis today.
The ADC noise spectra of current situation are show in Fig.15 ~ Fig.26.
knowledge, thoughts, etc.
Spectra in Fig.14 were measured with the ADC which had been used in ITMY.
Channel #2 had been assigned to ITMY_IP_ACC_H2 which was glitchy one (See also Fig.2 in klog#20636).
But it seemed to be healthy according to the measurements on the test bench reported as klog#20682.
Current test bench may not be enough for checking our problem
because the situation of power supply and ground connection was probably different so much from one in the mine.
When I measured ADC noise on the test bench, only one ADC board and one AA chassis were installed.
I plan to test ADCs and AAs which were brought back to Mozumi today with multi-ADC, and -AA installation.